The velocity v of the wave set-up by

plucking a stretched string is found to

depend on the tension T in the string, its

length l and its mass m, and is given by

 where x, y, z are unknown numbers and k is

a constant. Find the values of x, y, and z.


1
Expert's answer
2021-01-31T18:58:05-0500
"(\dim v)=(\dim t) ^
x
 (\dim l) ^
y
 (\dim m) ^
z"

"LT ^\n{\u22121}\n =(MLT ^{\n\u22122}\n ) ^\nx\n (L) ^\ny\n (M) ^\nz""\u22121=\u22122x\u2192x=\\frac{1}{2}\\\\1= \n\\frac{1}{2}\n\u200b\t\n +y\u2192y=\\frac{1}{2}\\\\0= \n\\frac{1}{2}\n\u200b\t\n +z\u2192z=-\\frac{1}{2}"